# Readout electronics for low occupancy High-Pressure Gas TPCs

### v1

@@ 4.2 Thermal profile
To achieve a reasonable gain in a gas TPC, a stable pressure must be maintained. The thermal profile of the system thus needs to be well-understood as in a closed vessel the temperature directly affects the pressure. Cooling is currently achieved in TOAD through thermal contact between the aluminium mounting plate of the PAT board and the OROC. The FECs are cooled through contact with the gas in the vessel. The temperature and the power output from each PAT board are recorded via the FEC DC/DC converter on each board, and the average values in are plotted in figure 16. The idle power consumption, with no power delivered to the FECs, for the PAT boards all connected to the same power bus ranges between 0.89-.
@@
Relying solely on the thermal coupling between the PAT boards and the OROC, an average change of temperature of on the DC/DC converter was observed as a result of the operation of the FECs. While the final temperature reached is within the operation threshold of the electronics, a large temperature change and use of the gas as the cooling medium can lead to a long stabilisation time for the pressure in the vessel as well as uncertain gas drift properties. As shown in figure 17, by fitting an exponential to the pressure profile, it takes for the pressure to have a rate of change of pressure . It is also worth noting that the heavy pressure vessel is in contact with the gas, increasing the time to heat up. Furthermore, the OROC is also in thermal contact with the TOAD vessel and the gas, further increasing the pressure stabilisation time and also exposing it to the high equilibrium temperature. Such a long stabilisation time prolongs the time taken to go from the detector being off to stable operations, increasing downtime.
@@ 5 Conclusion
Many readout systems developed for particle physics experiments target the high-occupancy environment of a hadron collider experiment. We have described a system designed instead for much lower occupancy detectors, such as those in neutrino physics experiments. The system leverages the lower buffering requirements to use far fewer FPGAs per channel and thereby significantly reduce the cost of the system.
@@
The system performed well, stable operation was seen throughout the test period which included multiple days of running at atmosphere and pressure as well as bench top testing. However, the passive cooling used in the test has been identified as an area that must be improved before final deployment. Nevertheless, the total heat load is well within the range that can be dealt with active cooling. The noise characteristics of the system were measured and found to be within the range that will allow us to read out the intended signals for almost all channels. We have identified poor connection due to the mounting of the FECs in the legacy connectors of the OROC as the primary item contributing to channels that are not readable. This is specific to the OROC used in TOAD, and is not expected to be a problem for future HPgTPC e.g. ND-GAr which will use different connectors. Notwithstanding these necessary minor modifications, we have shown that this system is able to fulfil the needs of a high-pressure gas TPC detector for a fraction of the price of systems used for LHC type detectors.

### v2

@@ 4.2 Thermal profile
To achieve a reasonable gain in a gas TPC, a stable pressure must be maintained. The thermal profile of the system thus needs to be well-understood as in a closed vessel the temperature directly affects the pressure. Cooling is currently achieved in TOAD through thermal contact between the aluminium mounting plate of the PAT board and the OROC. The FECs are cooled through contact with the gas in the vessel. The temperature and the power output from each PAT board are recorded via the FEC DC/DC converter on each board, and the average values in are plotted in figure˜15. The idle power consumption, with no power delivered to the FECs, for the PAT boards all connected to the same power bus, ranges between 0.89-. Additionally, the idle temperature in equilibrium with the environment has been found to be , which is well within the operational margin provided by AMD [18].
@@
Relying solely on the thermal coupling between the PAT boards and the OROC, an average change of temperature of on the DC/DC converter was observed as a result of the operation of the FECs. While the final temperature reached is within the operation threshold of the electronics, a large temperature change and use of the gas as the cooling medium can lead to a long stabilisation time for the pressure in the vessel as well as uncertain gas drift properties. As shown in figure˜16, by fitting an exponential to the pressure profile, it takes for the pressure to have a rate of change of pressure . It is also worth noting that the heavy pressure vessel is in contact with the gas, increasing the time to heat up. Furthermore, the OROC is also in thermal contact with the TOAD vessel and the gas, further increasing the pressure stabilisation time and also exposing it to the high equilibrium temperature. Such a long stabilisation time prolongs the time taken to go from the detector being off to stable operations, increasing downtime.
@@
Stable performance of the electronics was observed during the pressure operation. During the time-constrained test-beam period, we observed stability in pressure for the duration of multiple days, with no significant variation in the electronics performance observed. Similar stability was observed in atmosphere during the test-beam period, as well as throughout benchtop testing.
@@ 5 Conclusion
Many readout systems developed for particle physics experiments target the high-occupancy environment of a hadron collider experiment. Scaling systems such as readout electronics from sPHENIX or ALICE would be unnecessarily costly. Similarly, technology such as LArPix would require overclocking to meet the resolution requirements for ND-GAr. Commercial solutions, such as CAEN, would also be considerably costly. We have, instead, described a custom system designed for much lower occupancy detectors, such as those in neutrino physics experiments. The system leverages the lower buffering requirements to use far fewer FPGAs per channel and thereby significantly reduces the cost of the system. Even though it is a prototype system manufactured at a small scale, the readout for TOAD (including FECs, PAT cards and all the hardware to connect and power them) costs less than $ for channels. Scaling this up to the full DUNE Phase-II ND, ignoring likely economies of scale, gives a conservative cost estimate of $1.4 million.
@@
The system performed well, with stable operation seen throughout the test period, which included multiple days of running at atmosphere and pressure as well as benchtop testing. However, the passive cooling used in the test has been identified as an area that must be improved before final deployment. Nevertheless, the total heat load is well within the range that can be dealt with using active cooling. The noise characteristics of the system were measured and found to be within the range that will allow us to read out the intended signals for almost all channels. We have identified poor connection due to the mounting of the FECs in the legacy connectors of the OROC as the primary item contributing to channels that are not readable. This is specific to the OROC used in TOAD, and is not expected to be a problem for future HPgTPC, e.g. ND-GAr, which will use different connectors. Notwithstanding these necessary minor modifications, we have shown that this system is able to fulfill the needs of a high-pressure gas TPC detector for a fraction of the price of systems used for LHC type detectors.